# Sample Efficient Algorithms for Linear System Identification under Noisy Observations

### v1

@@ Abstract
In this paper, we focus on learning linear dynamical systems under noisy observations. In this setting, existing algorithms either yield biased parameter estimates, or suffer from large sample complexities. To address these issues, we adapt the instrumental variable method and the bias compensation method, originally proposed for error-in-variables models, to our setting and provide refined non-asymptotic analysis. Under mild conditions, our algorithms achieve superior sample complexities that match the best-known sample complexity for learning a fully observable system without observation noise.
@@ I Introduction
Dynamical systems play a central role in numerous real-world applications, including neuroscience [1, 2], climate [3], power grids [4], robotics [5, 6], and beyond. In many of the applications, the underlying system models are unknown, making it challenging to perform critical downstream tasks including state prediction [7], false detection [8], feedback control [9]. In response, researchers have been developing data-driven algorithms for learning the system models, which are often referred to as system identification (SYSID) algorithms [10, 11].
In the context of linear systems, recent research mainly focuses on two classes of systems, i.e. fully observable and partially observable systems. In the fully observable setting, the entire state vectors are observed without observation noises. To learn the linear system models, it suffices to directly perform least-squares on the state and input trajectories. Recent studies [12, 13, 14] have established non-asymptotic error bounds of order $\tilde{\mathcal{O}}(\sqrt{(n+m)/T})$ for an $n$-dimensional systems with $m$-dimensional inputs and $T$ data samples, which is optimal up to constants and logarithmic factors. In partially observable systems, the states are measured through a linear observer with observation noises. In this setting, the most widely used algorithms are the Ho-Kalman-based algorithms [15, 16]. These methods first estimate the input-to-output mapping of the system, which is composed of the so-called Markov parameters. The dynamical model, or the system matrices, is subsequently recovered from the Markov parameters. They achieve non-asymptotic error upper bounds of order $\tilde{\mathcal{O}}(\sqrt{n^{x}(m+n+d)/T})$ for systems with $n$-dimensional states, $m$-dimensional input, $d$-dimensional observations and $T$ data samples. Here $x$ is some positive integer depending on the specific algorithm.
@@
Here $x_{t}\in\mathbb{R}^{n},u_{t}\in\mathbb{R}^{m},w_{t}\in\mathbb{R}^{n}$ denote the state, input and process noise, respectively. $\widehat{x}_{t}\in\mathbb{R}^{n}$ and $\eta_{t}\in\mathbb{R}^{n}$ denote the state observation and observation noise. On one hand, directly performing least-squares on $(\widehat{x}_{t},u_{t})$ and $\widehat{x}_{t+1}$ to estimate $A$ and $B$, as in fully observable systems, leads to biased estimations. Because both $\widehat{x}_{t}$ and $\widehat{x}_{t+1}$ are contaminated by observation noises, the above problem becomes an error-in-variable (EIV) problem [17, 18] where the least-squares estimator is known to be biased and does not converge to the true parameter regardless of the sample size. On the other hand, we can also treat the system as a partially observable system with an identity observer matrix $C=I_{n}$ and apply Ho-Kalman-based algorithms. However, this approach leads to a suboptimal estimation error. This is because under the standard controllability assumption, Ho-Kalman-based algorithms need to estimate the Markov parameters up to order $n$, i.e. $\{A^{i}B\}_{i=0}^{n}$, from data. Although the estimation is unbiased, the algorithms need to learn approximately $mn/(m+n)$ times more parameters than the target system, leading to a sample complexity at least $mn/(m+n)$ times worse than the optimal.
In this paper, we propose sample-efficient algorithms to learn the system in Equation 1. Specifically, we adapt two classical methods from the EIV problems, i.e. the instrumental variable (IV) method and the bias-compensation (BC) method [17], to our setting. We provide non-asymptotic analysis of the proposed algorithms, establishing an error upper bound of order $\tilde{\mathcal{O}}(\sqrt{(m+n)/T})$ on the estimated system matrices under mild assumptions. Here $\tilde{\mathcal{O}}(\cdot)$ hides constants and logarithmic factors. Notably, this rate matches the error upper bound for learning a fully observable linear system without any observation noises [12]. Namely, our algorithms effectively mitigate the influence of the observation noises while only sacrifice some noise-related constants and logarithmic factors. Achieving this tight upper bound requires careful non-asymptotic analysis, which is fundamentally different from the asymptotic analysis in existing EIV literature [17, 18].
Notations. We use $\lesssim$ and $\gtrsim$ to hide absolute constants. We use $\tilde{\mathcal{O}}(\cdot)$ to hide all problem-related constants, absolute constants and logarithmic factors. For any positive integer $a$, let $[a]$ denote set $\{1,2,\cdots,a\}$. For matrix $M\in\mathbb{R}^{m\times n}$, we let $\sigma_{1}(M)\geq\sigma_{2}(M)\geq\cdots\geq\sigma_{\max\{m,n\}}(M)\geq 0$ denote its singular values and let $\left\|M\right\|$ denote its operator norm.
@@ II-A System Model & Learning Objective
[…]
@@ II-B Necessary Assumptions
Intuitively, it assumes the $A$ matrix is “stable” so that the system states do not blow up over a long time period.
@@ II-C The Naive Least-Squares Estimator
$\begin{split}E=\begin{bmatrix}A&B\end{bmatrix},\quad z_{t}=\begin{bmatrix}x_{t}\\ u_{t}\end{bmatrix},\quad\widehat{z}_{t}=\begin{bmatrix}\widehat{x}_{t}\\ u_{t}\end{bmatrix}.\end{split}$
@@
Notice that $\widehat{x}_{t+1}=E\left(\widehat{z}_{t}-\begin{bmatrix}\eta_{t}\\ 0\end{bmatrix}\right)+w_{t}+\eta_{t+1}=E\widehat{z}_{t}-A\eta_{t}+w_{t}+\eta_{t+1}$. Substituting back gives
@@
We note that the noise $-A\eta_{t}+w_{t}+\eta_{t+1}$ is correlated with the measurement $\widehat{z}_{t}^{\intercal}$, since both terms have $\eta_{t}$ as a component. To see how this correlation influence the performance of $\widehat{E}_{\text{LS}}$, we further decompose $\widehat{E}_{\text{LS}}$ as follows
@@
To eliminate this bias, we introduce two methods in the following sections, i.e. the instrumental vairable (IV) method and the bias compensation (BC) method, both adapted from previous research on single-input-single-output auto-regressive error-in-variable models [17].
@@ III-A The IV Estimator and Its Design Rationale
The IV method eliminates the bias in $\widehat{E}_{\text{LS}}$, or equivalently, eliminates the correlation between measurement $\widehat{z}_{t}$ and noise $-A\eta_{t}+w_{t}+\eta_{t+1}$, by replacing $\widehat{z}_{t}$ with an instrumental variable $\widehat{i}_{t}$ and modifying the estimator as follows
@@
On the high level, we require instrumental variables $\widehat{i}_{1:t}$ to be uncorrelated with $-A\eta_{t}+w_{t}+\eta_{t+1}$ so that the norm of covariance $\sum_{t=0}^{T-1}(-A\eta_{t}+w_{t}+\eta_{t+1})\widehat{i}_{t}^{\intercal}$ is small [^2] Asymptotically, for this norm to be small, it only requires $\widehat{i}_{t}$ to be uncorrelated with $-A\eta_{t}+w_{t}+\eta_{t+1}$. However, here we aim for a non-asymptotic upper bound and adopt a slightly strong requirement for technical reasons.. Moreover, we require $\widehat{i}_{t}$ to be well correlated with $\widehat{z}_{t}$ so that the norm of $\sum_{t=0}^{T-1}\widehat{z}_{t}\widehat{i}_{t}^{\intercal}$ is large. Combining the two requirements would lead to a small error term $\widehat{E}_{\text{IV}}-E$.
@@
$\begin{split}\widehat{i}_{t}=\begin{bmatrix}\widehat{x}_{t-1}\\ u_{t}\end{bmatrix}\in\mathbb{R}^{n+m}.\end{split}$ (5)
suffices. To see this, we first note that $\widehat{i}_{1:t}$ is uncorrelated with $-A\eta_{t}+w_{t}+\eta_{t+1}$. With this key fact, we can prove $\left\|\sum_{t=0}^{T-1}(-A\eta_{t}+w_{t}+\eta_{t+1})\widehat{i}_{t}^{\intercal}\right\|\leq\tilde{\mathcal{O}}\left(\sqrt{T}\right)$. Now to show $\widehat{i}_{t}$ strongly correlates with measurement $\widehat{z}_{t}$, we analyze their covariance $\mathbb{E}\widehat{z}_{t}\widehat{i}_{t}^{\intercal}=\mathbb{E}\begin{bmatrix}\widehat{x}_{t}\widehat{x}_{t-1}^{\intercal}&\widehat{x}_{t}u_{t}^{\intercal}\\ u_{t}\widehat{x}_{t-1}^{\intercal}&u_{t}u_{t}^{\intercal}\end{bmatrix}$. For the upper left block, we know $\mathbb{E}\left(\widehat{x}_{t}\widehat{x}_{t-1}^{\intercal}\right)=\mathbb{E}(Ax_{t-1}+Bu_{t-1}+w_{t-1}+\eta_{t})\left(x_{t-1}+\eta_{t-1}\right)^{\intercal}=A\mathbb{E}\left(x_{t-1}x_{t-1}^{\intercal}\right)$ because $u_{t-1},w_{t-1},\eta_{t-1}$ and $\eta_{t}$ are independently sampled from zero-mean distributions. Similarly, we conclude that the off diagonal blocks are 00. It then naturally follows that, $\mathbb{E}\widehat{z}_{t}\widehat{i}_{t}^{\intercal}=\begin{bmatrix}A\mathbb{E}\left(x_{t-1}x_{t-1}\right)&0\\ 0&\Sigma_{u}\end{bmatrix}$, which is invertible as long as $A$ is invertible. Utilizing this key insight, we can prove that $\sum_{t=0}^{T-1}\widehat{z}_{t}\widehat{i}_{t}^{\intercal}\succeq\tilde{\mathcal{O}}(T)I$ for any invertible $A$. Combining the above two inequalities, we can conclude $\left\|\widehat{E}_{\text{IV}}-E\right\|\leq\tilde{\mathcal{O}}(1/\sqrt{T})$, which decays to 00 as $T$ tends to infinity.
@@
III-B Theoretical Guarantee for The IV Method
Now we formally present the theoretical guarantee, along with additional assumptions on the target system $\mathcal{M}=(A,B,\Sigma_{w},\Sigma_{u},\Sigma_{\eta})$, for estimator $\widehat{E}_{\text{IV}}$ (LABEL:eq:iv_estimator).
@@ III-B Theoretical Guarantee for The IV Method
Suppose $A$ is invertible. Let $\phi_{A}>0$ denote its minimal singular value. ∎
@@
The error $\left\|\widehat{E}_{\text{IV}}-E\right\|$ decays to 00 as $T$ tends to infinity. Moreover, it scales on the order of $\tilde{\mathcal{O}}(\sqrt{(m+n)/T})$, matching the error bound for learning a fully observable system without observation noise [12]. Namely, our estimator effectively mitigates the influence of the observation noises while incurring only constant and logarithmic overhead.
@@ III-C Proofs for Theorem 1
[…]
$\begin{split}\sigma_{1}\left(\Sigma_{i}\right)\precsim{}&\frac{\psi\psi_{A}^{2}}{1-\rho_{A}^{2}}\log\left(\frac{4}{\delta}\right)nT\\ \sigma_{n}(\Sigma_{i})\geq{}&\min\left\{\frac{\phi_{R}^{2}}{16},\frac{3}{8}\right\}\phi_{u}T.\end{split}$
@@
[…]
@@
Therefore, comining Lemma 3, we have the following with probability at least $1-3\delta$
@@
Step 4: Error Upperbound $\mathit{\left\|\widehat{E}_{\text{IV}}-E\right\|}$. Finally, substituting Equations 16 and 15 back into Equation 9, we get the following with probability at least $1-11\delta$
@@ IV-A The BC Estimator and Its Design Rationale
$\begin{split}\widehat{E}_{\text{LS}}={}&E+\Delta_{1}-E\left(\sum_{t=0}^{T-1}\begin{bmatrix}\eta_{t}\\ 0\end{bmatrix}\begin{bmatrix}\eta_{t}\\ 0\end{bmatrix}^{\intercal}\right)\left(\sum_{t=0}^{T-1}\widehat{z}_{t}\widehat{z}_{t}^{\intercal}\right)^{-1}\\ ={}&E+\Delta_{1}-E\begin{bmatrix}\Sigma_{\eta}&0\\ 0&0\end{bmatrix}\left(\frac{1}{T}\sum_{t=0}^{T-1}\widehat{z}_{t}\widehat{z}_{t}^{\intercal}\right)^{-1}\\ {}&+E\begin{bmatrix}\Sigma_{\eta}-\frac{1}{T}\sum_{t=0}^{T-1}\eta_{t}\eta_{t}^{\intercal}&0\\ 0&0\end{bmatrix}\left(\frac{1}{T}\sum_{t=0}^{T-1}\widehat{z}_{t}\widehat{z}_{t}^{\intercal}\right)^{-1}\\ \end{split}$

### v2

# Error-In-Variables Methods for Efficient System Identification with Finite-Sample Guarantees

@@ Abstract
This paper addresses the problem of learning linear dynamical systems from noisy observations. In this setting, existing algorithms either yield biased parameter estimates or have large sample complexities. We resolve these issues by adapting the instrumental variable method and the bias compensation method, originally proposed for error-in-variables models, to our setting. We provide refined non-asymptotic analysis for both methods. Under mild conditions, our algorithms achieve superior sample complexities that match the best-known sample complexity for learning a fully observable system without observation noise.
@@ I Introduction
Dynamical systems play a central role in numerous real-world applications, spanning neuroscience [1, 2], climate [3], power grids [4], robotics [5, 6], and beyond. In many of these domains, the underlying system models are unknown, making it challenging to perform critical downstream tasks including state prediction [7], false detection [8], and feedback control [9]. In response, researchers have been developing data-driven algorithms for learning the system models, which are often referred to as system identification (SYSID) algorithms [10, 11].
In the context of linear systems, recent research primarily focuses on two classes of systems, i.e. fully observable and partially observable systems. In the fully observable setting, the entire state vectors are observed without observation noises. To learn the linear system models, it suffices to perform least-squares on the state and input trajectories. Recent studies [12, 13, 14] have established non-asymptotic error bounds of order $\tilde{\mathcal{O}}(\sqrt{(n+m)/T})$ for an $n$-dimensional systems with $m$-dimensional inputs and $T$ data samples, which are optimal up to constants and logarithmic factors. In partially observable systems, the states are measured through a linear observer with observation noises. In this setting, the most widely used algorithms are Ho-Kalman-based algorithms [15, 16]. These methods first estimate the input-to-output mapping of the system, which is composed of the so-called Markov parameters. The dynamical model, or the system matrices, is subsequently recovered from the Markov parameters. They achieve non-asymptotic error upper bounds of order $\tilde{\mathcal{O}}(\sqrt{n^{x}(m+n+d)/T})$ for systems with $n$-dimensional states, $m$-dimensional input, $d$-dimensional observations and $T$ data samples. Here $x$ is some positive integer depending on the specific algorithm.
@@
Here $x_{t}\in\mathbb{R}^{n},u_{t}\in\mathbb{R}^{m},w_{t}\in\mathbb{R}^{n}$ denote the state, input and process noise, respectively. $\widehat{x}_{t}\in\mathbb{R}^{n}$ and $\eta_{t}\in\mathbb{R}^{n}$ denote the state observation and observation noise. On one hand, directly applying least-squares on $(\widehat{x}_{t},u_{t})$ and $\widehat{x}_{t+1}$, as in the fully observable case, leads to biased estimations of matrices $A$ and $B$. Because both $\widehat{x}_{t}$ and $\widehat{x}_{t+1}$ are contaminated by observation noises, the above problem becomes an error-in-variable (EIV) problem [17, 18] where the least-squares estimator is known to be biased and does not converge to the true parameter regardless of the number of samples. On the other hand, we can treat the system as a partially observable system with an identity observer matrix $C=I_{n}$ and apply Ho-Kalman-based algorithms. However, this approach leads to a suboptimal estimation error. This is because under the standard controllability assumption, Ho-Kalman-based algorithms need to estimate the Markov parameters up to order $n$, i.e. $\{A^{i}B\}_{i=0}^{n}$, from data. Although the estimation is unbiased, the algorithms need to learn approximately $mn/(m+n)$ times more parameters than necessary, thereby inflating the sample complexity by at least a factor of $mn/(m+n)$.
[…]
@@ II-A System Model & Learning Objective
We study the linear dynamical system in Equation 1 with $n$-dimensional states and $m$-dimensional inputs. For simplicity, we assume $x_{0}=0$. We also assume $u_{t}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,\Sigma_{u})$, $w_{t}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,\Sigma_{w})$, $\eta_{t}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,\Sigma_{\eta})$ [^1] Our analysis can be easily adapted to inputs and noises with subGaussian distributions.. We denote the above system, along with noises in the inputs and observation noises, by $\mathcal{M}=(A,B,\Sigma_{w},\Sigma_{u},\Sigma_{\eta})$.
@@ II-B Necessary Assumptions
Intuitively, it assumes the $A$ matrix is “stable” so that the system states do not blow up over a long horizon.
@@ II-C The Naive Least-Squares Estimator
[…]
@@
Notice that $\widehat{x}_{t+1}=E\left(\widehat{z}_{t}-\begin{bmatrix}\eta_{t}\\ 0\end{bmatrix}\right)+w_{t}+\eta_{t+1}=E\widehat{z}_{t}-A\eta_{t}+w_{t}+\eta_{t+1}$. Substituting this into Equation 2 gives
@@
We note that the noise $A\eta_{t}+w_{t}+\eta_{t+1}$ is correlated with the measurement $\widehat{z}_{t}^{\intercal}$, since both terms have $\eta_{t}$ as a component. To see how this correlation influences the performance of $\widehat{E}_{\text{LS}}$, we further decompose $\widehat{E}_{\text{LS}}$ as follows
@@
To eliminate this bias, we introduce two methods in the following sections, i.e. the instrumental variable (IV) method and the bias compensation (BC) method, both adapted from previous research on single-input-single-output auto-regressive error-in-variable models [17].
@@ III-A The IV Estimator and Its Design Rationale
The IV method eliminates the bias in $\widehat{E}_{\text{LS}}$, or equivalently, eliminates the correlation between measurement $\widehat{z}_{t}$ and noise $A\eta_{t}+w_{t}+\eta_{t+1}$, by replacing $\widehat{z}_{t}$ with an instrumental variable $\widehat{i}_{t}$ and modifying the estimator as follows
@@
On the high level, we require instrumental variables $\widehat{i}_{1:t}$ to be uncorrelated with $A\eta_{t}+w_{t}+\eta_{t+1}$ so that the norm of covariance $\sum_{t=0}^{T-1}(-A\eta_{t}+w_{t}+\eta_{t+1})\widehat{i}_{t}^{\intercal}$ is small [^2] Asymptotically, for this norm to be small, it only requires $\widehat{i}_{t}$ to be uncorrelated with $A\eta_{t}+w_{t}+\eta_{t+1}$. However, here we aim for a non-asymptotic upper bound and adopt a slightly strong requirement for technical reasons.. Moreover, we require $\widehat{i}_{t}$ to be well correlated with $\widehat{z}_{t}$ so that the norm of $\sum_{t=0}^{T-1}\widehat{z}_{t}\widehat{i}_{t}^{\intercal}$ is large. Combining the two requirements would lead to a small error term $\widehat{E}_{\text{IV}}-E$.
@@
$\begin{split}\widehat{i}_{t}=\begin{bmatrix}\widehat{x}_{t-1}\\ u_{t}\end{bmatrix}\in\mathbb{R}^{n+m}\end{split}$ (5)
suffices. To see this, we first note that $\widehat{i}_{1:t}$ is uncorrelated with $A\eta_{t}+w_{t}+\eta_{t+1}$. We can then prove $\left\|\sum_{t=0}^{T-1}(-A\eta_{t}+w_{t}+\eta_{t+1})\widehat{i}_{t}^{\intercal}\right\|\leq\tilde{\mathcal{O}}\left(\sqrt{T}\right)$. Secondly, we show $\widehat{i}_{t}$ strongly correlates with measurement $\widehat{z}_{t}$ by analyzing their covariance $\mathbb{E}\widehat{z}_{t}\widehat{i}_{t}^{\intercal}=\mathbb{E}\begin{bmatrix}\widehat{x}_{t}\widehat{x}_{t-1}^{\intercal}&\widehat{x}_{t}u_{t}^{\intercal}\\ u_{t}\widehat{x}_{t-1}^{\intercal}&u_{t}u_{t}^{\intercal}\end{bmatrix}$. For the upper left block, we know $\mathbb{E}\left(\widehat{x}_{t}\widehat{x}_{t-1}^{\intercal}\right)=\mathbb{E}(Ax_{t-1}+Bu_{t-1}+w_{t-1}+\eta_{t})\left(x_{t-1}+\eta_{t-1}\right)^{\intercal}=A\mathbb{E}\left(x_{t-1}x_{t-1}^{\intercal}\right)$ because $u_{t-1},w_{t-1},\eta_{t-1}$ and $\eta_{t}$ are independently sampled from zero-mean distributions. Similarly, we conclude that the off diagonal blocks are 0. It then naturally follows that, $\mathbb{E}\widehat{z}_{t}\widehat{i}_{t}^{\intercal}=\begin{bmatrix}A\mathbb{E}\left(x_{t-1}x_{t-1}\right)&0\\ 0&\Sigma_{u}\end{bmatrix}$, which is invertible as long as $A$ is invertible. Utilizing this key insight, we can prove that $\sum_{t=0}^{T-1}\widehat{z}_{t}\widehat{i}_{t}^{\intercal}\succeq\tilde{\mathcal{O}}(T)I$ for any invertible $A$. Combining the above two inequalities, we can conclude $\left\|\widehat{E}_{\text{IV}}-E\right\|\leq\tilde{\mathcal{O}}(1/\sqrt{T})$, which decays to 0 as $T$ tends to infinity.
@@
III-B Theoretical Guarantee for The IV Estimator
Now we formally present the finite sample error bound on the IV estimator (LABEL:{eq:iv_estimator}), along with additional assumptions on the target system $\mathcal{M}=(A,B,\Sigma_{w},\Sigma_{u},\Sigma_{\eta})$.
@@ III-B Theoretical Guarantee for The IV Estimator
AA is invertible. Let $\phi_{A}>0$ denote its minimal singular value. ∎
@@
The error $\left\|\widehat{E}_{\text{IV}}-E\right\|$ decays to 0 as $T$ tends to infinity. Moreover, it scales on the order of $\tilde{\mathcal{O}}(\sqrt{(m+n)/T})$, matching the error bound to learn a fully observable system without observation noise [12]. Thus, our estimator effectively mitigates the influence of the observation noises while incurring only constant and logarithmic overhead.
@@ III-C Proofs for Theorem 1
To bound the second term in Equation 10, we notice that $\eta_{t+1}+w_{t}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,\Sigma_{\eta}+\Sigma_{w})$ and is independent of $i_{1:t}$, i.e. independent of $u_{1:t}$ and $x_{0:t-1}$. Moreover, by Lemma 1, with probability at least $1-\delta$
$\begin{split}\sigma_{1}\left(\Sigma_{i}\right)\precsim{}&\frac{\psi\psi_{A}^{2}}{1-\rho_{A}^{2}}\log\left(\frac{4}{\delta}\right)nT\\ \sigma_{n}(\Sigma_{i})\geq{}&\min\left\{\frac{\phi_{R}^{2}}{16},\frac{3}{8}\right\}\phi_{u}T.\end{split}$ (13)
@@
Similarly, for the third term in Equation 10, we note that $\eta_{t+1}+w_{t}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,\Sigma_{\eta}+\Sigma_{w})$ and is independent of $\eta_{0:t-1}$. Moreover, we define $\Sigma_{\tilde{\eta}}\coloneqq\sum_{t=0}^{T-2}\begin{bmatrix}\eta_{t}\\ 0\end{bmatrix}\begin{bmatrix}\eta_{t}\\ 0\end{bmatrix}^{\intercal}$ and get the following by a standard covariance concentration result (e.g. Equation (45) of [21]) with probability at least $1-\delta$ for $T$ satisfying Equation 7
@@
Therefore, combining Equation 13, we have the following with probability at least $1-3\delta$
@@
Step 4: Error Upperbound $\mathit{\left\|\widehat{E}_{\text{IV}}-E\right\|}$. Finally, substituting Equations 17 and 16 back into Equation 9, we get the following with probability at least $1-11\delta$
@@ IV-A The BC Estimator and Its Design Rationale
[…]